The velocity of waves on a string is 99 m/s. If the frequency of standing waves is 470 Hz, how far apart are two adjacent nodes?

_______ m
Answer #1

The wavel ength of standing waves is

$$ \lambda=\frac{v}{f}=\frac{99 \mathrm{~m} / \mathrm{s}}{470 \mathrm{~Hz}}=0.2106 \mathrm{r} $$

The distance between two adjacent nodes is

$$ d=\frac{\lambda}{2}=\frac{0.2106 \mathrm{~m}}{2}=0.105 \mathrm{~m} $$